# Speech emotion recognition from voice messages recorded in the wild

### v1

@@ II Related work
Finally, research on evaluating model fairness is limited, particularly in the context of pre-trained models [27]. Some previous works have investigated gender-based fairness in Spanish databases, revealing differences in model performance between females and males. In [38], the authors reported a female accuracy of 90.9% and a male accuracy of 89.4%, yet the database only included one male and one female speaker. Conversely, [21] and [42] showed a different pattern, with males achieving higher accuracy than females. Specifically, [21] presented 89.49% for females and 93.90% for males (with 4 female and 4 male speakers), while [42] obtained 91.8% for females and 97.7% for males using pre-trained models (with 33 male and 24 female speakers). The same trend has been found in some works on SER for English data. In [27], they studied the fairness of SER systems using pre-trained models and observed a reduction of 0.234 in CCC for arousal among females as opposed to males on MSP-Podcast. In [11], they found that pre-trained models tend to exhibit greater fairness in predicting arousal and dominance than in valence. Notably, for valence, the majority of models showed higher CCC for females than for males. Nevertheless, overall the speech representations obtained with pre-trained models seem to be invariant to domain, speaker, and gender. Additionally, the authors in [11] also explored fairness across individual speakers and found that different pre-trained models show overall consensus on categorizing speakers as ‘good’ or ‘bad’, obtaining lower CCC values for some individuals in the latter group. These findings underline the importance of incorporating fairness assessment in future research.
III Speech emotion databases
This research used three databases: EMOVOME, IEMOCAP and RAVDESS, detailed below. A comparison of the main features is presented in Table II, indicating the type of speech (acted, elicited or natural), the language, the labeled emotions categories or dimensions, the number of samples (N), the number of speakers (Spk) with the gender distribution (males/females), the mean number of samples per speaker (N/Spk) and the sample duration (mean and range).
@@ III Speech emotion databases
TABLE II: Comparison of the emotion databases used in this work: EMOVOME, RAVDESS and IEMOCAP.
Database Type Language Dimensions Categories N Spk (M/F) N/Spk Sample duration
@@
EMOVOME Natural Spanish Valence, arousal Big six + neutral 999 100 (50/50) 10 17.59s 1-60s
@@
Elicited,
acted
English
Valence, arousal,
dominance
Big six + neutral,
frustration, excited
10039 10 (5/5) 1004 4.59s 1-35s
RAVDESS Acted English - Big six + neutral, calm 1440 24 (12/12) 60 3.70s 3-5s
III-A Emotional Voice Messages database (EMOVOME)
The Emotional Voice Messages (EMOVOME) database [29] was created from scratch for this work to obtain emotional speeches in real-world conditions. It contains 999 audio messages collected from real WhatsApp conversations of 100 Spanish speakers (50 female, 50 male). Voice messages were produced in-the-wild conditions before participants were recruited, avoiding any conscious bias due to the laboratory environment. Samples were labeled by two clinical psychologists (considered experts in the task of recognizing emotions) and three other annotators (deemed non-experts) in terms of valence and arousal using a 5-point scale. The experts labeled half of the audios each, so henceforth it will be considered a single expert label. The labels were processed to obtain three categories for arousal (high, neutral and low) and valence (positive, neutral and negative). They were finally aggregated by majority voting to obtain a single label per audio, giving priority to the expert’s label in case of a tie. The expert also labeled the audios in the following categories: happy, angry, fearful, sad, surprise, disgust and neutral. In this work, only the four most frequent emotions were used for classification, i.e., happy, angry, neutral and surprise. Details of the experimental procedure used for data collection and labeling are described in [29]. To the best of our knowledge, EMOVOME is the first collection of spontaneous emotions from real voice messages.
III-B Interactive Emotional Dyadic Motion Capture database (IEMOCAP)
@@
The Interactive Emotional Dyadic Motion Capture (IEMOCAP) database [30] stands out as a benchmark dataset for studying emotional expression and communication. It contains dyadic interactions of 10 English-speaking actors engaged in scripted and improvised dialogues designed to elicit different emotions. In this work, only the audio files corresponding to the segmentation of the conversations into utterances were used to create the SER models. This set includes 10039 utterances annotated into 9 emotion categories and 3 dimensions.
To align the classification results with the EMOVOME database, valence and arousal scores were stratified into three categories. Samples with a score below 2.5 were considered negative/low valence/arousal. Those samples with scores between 2.5 and 3.5 were considered neutral valence/arousal. Finally, we considered samples as positive/high valence/arousal if the score was greater than 3.5. Moreover, to facilitate comparison with the literature, only the four most frequent emotions were used for the classification into categories, i.e., angry, neutral, sad, happy and excited (the latter two were merged, following previous studies).
III-C Ryerson Audio-Visual Database of Emotional Speech and Song (RAVDESS)
The Ryerson Audio-Visual Database of Emotional Speech and Song (RAVDESS) [31] is a multimodal database that contains speech and song recordings expressing different emotional states. In this investigation, the voice-only data from RAVDESS was used. In this set, 24 actors pronounce two phrases (“Kids are talking by the door” and “Dogs are sitting by the door”) with different emotional intentions in a recording studio. Actors vocalized each statement in normal and strong emotional intensity for each of the eight emotions (except for neutral), resulting in 1440 samples.
@@ III-C Ryerson Audio-Visual Database of Emotional Speech and Song (RAVDESS)
To compare the classification results with the EMOVOME database, the emotion labels were transformed from the discrete emotional model to the dimensional emotional model, as also studied in [49]. For this purpose, the eight emotion categories were converted into valence and arousal dimensions following their distribution in Russell’s circumplex model of affect [5] (see Fig. 1). The label conversion in the valence dimension was the following: negative (sad, angry, fearful, disgust), neutral (neutral, surprised) and positive (happy, calm). Likewise, the label conversion for the arousal dimension was: low (calm, sad), neutral (neutral, disgust) and high (happy, angry, fearful, surprised).
IV Methods
@@
Three approaches were implemented to create the SER models. First, a standard feature set with classical machine learning algorithms was used as a baseline (Section IV-A). Then, pre-trained models were used as feature extractors, followed by a linear layer for classification (Section IV-B). Finally, a combination of the pre-trained models and the standard features was analyzed (Section IV-C).
To evaluate the three methods, the speech databases were divided into 80% for development and 20% for testing using a speaker-independent approach. Details of the label distribution in each partition are provided in Table III. To facilitate reproducibility and comparison of results, for test, we used: the test set in [29] for EMOVOME; session 5 for IEMOCAP; and “fold 0” proposed in [50] for RAVDESS. Additionally, parameter tuning also employed a SI cross-validation scheme, utilizing the StratifiedGroupKFold from Scikit-learn [51], with 4 folds for IEMOCAP and 5 folds for EMOVOME and RAVDESS. Folds are formed by grouping all audio samples from the same subject into a single fold, ensuring a roughly equal distribution of labels within each fold. All methods were assessed using two evaluation metrics. Accuracy, also called weighted accuracy (WA), evaluates the ratio of correctly predicted class samples to the total samples. WA is suitable for balanced datasets but less so for imbalanced ones, as it gives more weight to classes with more samples. In SER literature, unweighted accuracy (UA) is often preferred, representing the average of individual class accuracies. The Python libraries LibROSA, Scikit-learn, TensorFlow and Keras were used to implement the models.
TABLE III: Data distribution in training and test partitions of EMOVOME, IEMOCAP and RAVDESS.
Database Spk (M/F) Valence Arousal Categories
Negative Neutral Positive Low Neutral High hap ang neu sur sad fea dis cal
Training
EMOVOME - E 80 (40/40) 251 247 309 265 347 195 264 157 184 105 - - - -
EMOVOME - N 80 (40/40) 219 363 219 76 244 411 - - - - - - - -
EMOVOME - C 80 (40/40) 241 305 261 148 328 331 - - - - - - - -
IEMOCAP 8 (4/4) 2622 3725 1522 1023 5405 1441 1194 933 1324 - 839 - - -
RAVDESS 19 (10/9) 608 228 304 304 228 608 152 152 76 152 152 152 152 152
Test
EMOVOME - E 20 (10/10) 59 46 87 34 76 82 78 42 41 13 - - - -
EMOVOME - N 20 (10/10) 56 70 66 13 51 128 - - - - - - - -
EMOVOME - C 20 (10/10) 57 62 73 13 70 109 - - - - - - - -
IEMOCAP 2 (1/1) 799 883 488 258 1503 409 442 170 384 - 245 - - -
RAVDESS 5 (2/3) 160 60 80 80 60 160 40 40 20 40 40 40 40 40
IV-A Baseline: eGeMAPS and machine learning
@@
As a first approach, we repeated the baseline method proposed in [29] using long-term acoustic features and machine learning algorithms. First, audio samples in EMOVOME were resampled to 44.1 kHz, while for RAVDESS and IEMOCAP, the sample rate was kept at 48 kHz and 16 kHz, respectively. Next, the extended Geneva Minimalistic Acoustic Parameter Set (eGeMAPS) [52] was extracted using the openSMILE toolkit [53]. The 88 features per audio were normalized by subtracting the mean and dividing by the standard deviation of the development samples. For feature selection, high-correlated features ($p>0.95$) were first eliminated using Pearson’s correlation matrix, and a filter method was then used to select 25%, 50% or 75% of the features based on the highest ANOVA F-values. Finally, Support Vector Machine (SVM) and K-Nearest Neighbours (KNN) models were fitted on the development set according to the cross-validation scheme indicated above, tuning the combination of hyperparameters detailed in [29]. The chosen combination of features and hyperparameters was used to train a model on the entire development set, followed by an evaluation of the test set.
IV-B Pre-trained model embeddings
@@
We evaluated several pre-trained models with different architectures, pre-training methodologies and pre-training audio data, indicated in Table IV. First, several variations of the widely used Wav2vec 2.0 model were selected. Considering that EMOVOME is in Spanish, and we wanted to compare it with other two databases in English, we selected models pre-trained using a multilingual approach: facebook/wav2vec2-xls-r-300m (w2v2-xlsr-128) [54] pre-trained on 436k hours of audios in 128 languages, and facebook/wav2vec2-large-xlsr-53 (w2v2-xlsr-53) [55], pre-trained on 56k hours of audios in 53 languages. Both models included Spanish in the pre-training data. Despite prior research suggesting that fine-tuning models for automatic speech recognition do not help with speech emotion recognition [56, 12], we decided to include a fine-tuned model for Spanish ASR. This decision aimed to explore whether this approach outperforms the utilization of a pre-trained model in a different language since [56] and [12] used pre-trained models in the same language as the database tested (English). Therefore, we also used the fine-tuned version facebook/wav2vec2-large-xlsr-53-spanish (w2v2-xlsr-53-spa) [55] to obtain the embeddings. We also included a model pre-trained using noisy audios, the model facebook/wav2vec2-large-robust (w2v2-L-robust) [57], as it may be useful for the EMOVOME database despite being in English. As a widely-used alternative to Wav2Vec2, we used a HuBERT model, particularly the large version facebook/hubert-large-ll60k (hubert-L) [10]. Additionally, recent studies [45, 58, 59] have indicated that including information about the speaker is helpful for speech emotion recognition. Therefore, following these investigations, two more models were selected: Microsoft’s UniSpeech-SAT-Large (unispeech-L) [60] and a Statistics Pooling Time Delay Neural Network to obtain x-vector embeddings [61]. The former is a model pre-trained using multitask learning, including also the speaker identity during training. The latter provides a speaker embedding learning during a speaker verification task. All the pre-trained models are available in HuggingFace. The x-vector also requires the Speechbrain toolkit [62]. All of them require the input audio to be resampled to 16 kHz.
TABLE IV: Pre-trained models used in this study.
Model name HuggingFace name
w2v2-xlsr-128 facebook/wav2vec2-xls-r-300m
w2v2-xlsr-53 facebook/wav2vec2-large-xlsr-53
w2v2-xlsr-53-spa facebook/wav2vec2-large-xlsr-53-spanish
w2v2-L-robust facebook/wav2vec2-large-robust
hubert-L facebook/hubert-large-ll60k
unispeech-L microsoft/unispeech-sat-large
x-vector speechbrain/spkrec-xvect-voxceleb
@@
IV-C Pre-trained model embeddings and eGeMAPS

### v3

@@ II-C Fairness evaluation
Finally, research on evaluating model fairness is limited, particularly in the context of pre-trained models [46]. Some previous works have investigated gender-based fairness in Spanish datasets, revealing differences in model performance between females and males. In [33], the authors reported a female accuracy of 90.9% and a male accuracy of 89.4%, yet the dataset only included one male and one female speaker. Conversely, [30] and [38] showed a different pattern, with males achieving higher accuracy than females. Specifically, [30] presented 89.49% for females and 93.90% for males (with 4 female and 4 male speakers), while [38] obtained 91.8% for females and 97.7% for males using pre-trained models (with 33 male and 24 female speakers). The same trend has been found in some works on SER for English data. In [46], they studied the fairness of SER systems using pre-trained models and observed a reduction of 0.234 in CCC for arousal among females as opposed to males on MSP-Podcast. In [40], they found that pre-trained models tend to exhibit greater fairness in predicting arousal and dominance than in valence. Notably, the majority of models showed higher CCC for females than for males for valence. Nevertheless, overall the speech representations obtained with pre-trained models seem to be invariant to domain, speaker, and gender. Additionally, the authors in [40] also explored fairness across individual speakers and found that different pre-trained models show overall consensus on categorizing speakers as ‘good’ or ‘bad’, obtaining lower CCC values for some individuals in the latter group. These findings underline the importance of incorporating fairness assessment in future research.
III Emotional Voice Messages (EMOVOME)
The Emotional Voice Messages (EMOVOME) dataset was created from scratch for this study is the first collection of spontaneous emotions from real voice messages. Figure 1 presents a diagram illustrating the steps followed to create the EMOVOME dataset, detailed in the following subsections. The EMOVOME dataset is publicly available at the following Zenodo repository https://zenodo.org/records/10694370. Access to the raw audio files requires completing and signing an agreement form to ensure the speakers’ privacy.
Figure 1: Overview of the methodology.
III-A Participants
The dataset includes 100 Spanish speakers (50% females, 50% males) equally distributed across four age ranges (18-25, 26-35, 36-45, and 46-55 years old), with no self-reported speech disorder, and who regularly send audio messages to their contacts. Additional demographic information can be found in the Appendix. Participants were sourced via a recruitment agency, which identified suitable individuals from their dataset based on the specified criteria and invited them to participate in the study for a monetary compensation of 25€. All methods and experimental protocols were performed in accordance with the regulations of the local ethics committee of the Universitat Politècnica de València (No. P02_04_06_20).
III-B Data collection
The sample collection was conducted through a web-based application designed for the study. Participants used their computer to complete the study, following on-screen instructions. Initially, they read the study protocol, provided informed consent, and completed a sociodemographic questionnaire. They then recorded an audio reading the short text in the Appendix. They also completed the NEO Five Factor Inventory (NEO-FFI) [47], a 60-item questionnaire assessing personality traits. Next, participants were asked to upload 12 voice messages they had previously sent to other contacts, with a third of them being positive, neutral, and negative to ensure a balanced sample in terms of valence. Valence was chosen over specific emotion categories to simplify the task, as obtaining multiple samples for each distinct emotion can be challenging. Platform instructions clarified the concepts of positive, neutral, and negative valence, providing examples for each. Voice messages were encoded using the Ogg Vorbis format.
III-C Quality check
Voice messages were generated in real-life settings before participants were recruited, minimizing any bias associated with laboratory and staged environments. Recordings took place in participants’ natural surroundings, ranging from quiet rooms at home to busy streets. To ensure high quality data, all recordings underwent a manual screening process. Audio files with critical background noise conditions (such as microphone malfunctions, background music or TV sounds) were identified and rejected. Participants were instructed to upload low-noise recordings, so if an audio was not considered suitable for the study, they were asked to send a new audio. We also excluded recordings that seemed to be recorded specifically for the study, as some participants may have attempted to fraudulently create these recordings after recruitment, despite instructions to upload voice messages from prior conversations. As a result, we received 1605 audio recordings, but 574 were not included in EMOVOME for this reason, resulting in 1031 audios. Finally, 32 audios exceeding 60 seconds were rejected to obtain a homogeneous sample concerning duration. Therefore, a total of 999 audios were considered in the final dataset (total duration of 293min), plus 100 audios reading the text.
III-D Data labeling
The emotional content of the audios was labeled along two dimensions: valence, i.e., the degree to which an emotion is perceived as positive or negative; and arousal, i.e., how strongly the emotion is felt. Three non-experts and two expert evaluators were recruited for the task. The latter consisted of two clinical psychologists who rated half of the audios each. They are considered experts due to their professional training in recognizing and interpreting emotions. The demographic information of the annotators is indicated in the Appendix. Using the Self-Assessment Manikin (SAM) procedure [48], the evaluators rated each audio’s valence and arousal on a 5-point pictorial scale. Additionally, the experts provided an extra label for 7 emotion categories: the six basic emotions (happiness, disgust, anger, surprise, fear and sadness) defined by Ekman [1] plus a neutral category.
The valence and arousal ratings of the non-experts and expert evaluators present a V-shaped relation, that is, arousal increases with positive or negative valence (see the Appendix). To compare the ratings among evaluators, three categories for valence (positive, neutral and negative) and three for arousal (high, neutral and low) were defined by grouping negative and positive scores in the scale [-2, 2] and keeping 0 as neutral. The agreement between evaluators is presented in Table III using Cohen’s kappa score, for both valence (V) and arousal (A). For valence, the kappa score goes from 0.614 to 0.754, indicating substantial agreement among annotators. In contrast, the agreement decreased considerably for the arousal dimension, with kappa values ranging from 0.075 to a maximum of 0.407. These results suggest valence recognition was easier for raters than arousal recognition.
TABLE III: Pair-wise Cohen Kappa scores for non-experts (NE) and experts (E) pairs, for both valence (V) and arousal (A)
NE 1 NE 2 NE 3 E
NE 1 -
V = 0.700
A = 0.142
V = 0.754
A = 0.407
V = 0.614
A = 0.275
NE 2
V = 0.700
A = 0.142
V = 0.732
A = 0.235
V = 0.646
A = 0.075
NE 3
V = 0.754
A = 0.407
V = 0.732
A = 0.235
V = 0.649
A = 0.241
E
V = 0.614
A = 0.275
V = 0.646
A = 0.075
V = 0.649
A = 0.241
The non-experts and expert evaluations were combined to obtain a final audio label in terms of valence and arousal using the mode, i.e. the most frequent category among the four assessments. In instances of a two-way tie, we prioritized the expert’s ratings as the final label for the audio. The distribution of the audio labels is shown in Figure 2. Concerning valence, it is roughly balanced, with neutral being the majority class ($\sim$37%) and negative the minority class ($\sim$30%). Regarding arousal, the distribution is clearly unbalanced, with $\sim$44high, $\sim$40neutral and $\sim$16low arousal. Considering the combination of valence and arousal labels, the majority clusters are [positive valence, high arousal], [neutral valence, neutral arousal] and [negative valence, high arousal], with 244, 224 and 152 audios, respectively. The remaining groups have less than 100 audios, with a minimum of 12 audios in the cluster [positive valence, low arousal].
Figure 2: Distribution of audio samples based on their arousal and valence labels. The circle area is proportional to the number of samples in each group.
Additionally, the expert annotators provided an extra label corresponding to 7 categories of emotions. The classes are distributed as follows: happiness (342), disgust (8), anger (199), surprise (118), fear (35), sadness (72) and neutral (225). Figure 3 shows the correspondence between categories and the valence and arousal dimensions provided by the experts. In this work, only the four most frequent emotions were used for classification, i.e., happiness, anger, neutral and surprise.
Figure 3: Valence and arousal ratings for each emotion. The diameter represents the number of samples using a logarithmic scale
III-E Transcription
In addition to the speech files, audio transcriptions have been included in the dataset to offer an extra modality for analysis. The voice messages were transcribed using Amazon Transcribe and subsequently manually corrected to ensure accuracy and reliability.
IV Reference datasets: IEMOCAP and RAVDESS
In order to compare the novel natural dataset EMOVOME with the current state of the art, we selected two of the most used datasets: IEMOCAP and RAVDESS. A comparison of the main features is presented in Table IV.
@@ IV Reference datasets: IEMOCAP and RAVDESS
TABLE IV: Comparison of the emotion datasets used in this work: EMOVOME, IEMOCAP and RAVDESS.
Dataset Expression authenticity Recording environment Language Emotions Raters Samples Speakers (Male/Female) Samples/ speaker Sample duration
@@
EMOVOME Natural Real-life Spanish
Continuous
Discrete (7)
4 999 100 (50/50) 10 17.59s 1-60s
@@
Elicited
and acted
Lab English
Continuous
Discrete (9)
319 10039 10 (5/5) 1004 4.59s 1-35s
RAVDESS Acted Lab English Discrete (8) - 1440 24 (12/12) 60 3.70s 3-5s
@@
IEMOCAP [14] stands out as a benchmark dataset for studying emotional expression and communication. It contains dyadic interactions of 10 English-speaking actors engaged in scripted and improvised dialogues designed to elicit different emotions. In this work, only the audio files corresponding to the segmentation of the conversations into utterances were used to create the SER models. This set includes 10039 utterances annotated into 9 emotion categories and 3 dimensions. To align the classification results with the EMOVOME dataset, valence and arousal scores were stratified into three categories. Samples scoring below 2.5 were deemed negative/low, those between 2.5 and 3.5 were neutral, and those above 3.5 were positive/high. To facilitate comparison with the literature, only the four most frequent emotions were used for the classification into categories, i.e., angry, neutral, sad, happy and excited (the latter two were merged, following previous studies).
@@
RAVDESS [15] is a multimodal dataset with speech and song recordings expressing different emotions. In this study, only the voice data was used, where 24 actors recorded two phrases with normal and strong emotional intensity for each of eight emotions (excluding neutral), totaling 1,440 samples. To compare the results with EMOVOME, the emotion labels were transformed from the discrete model to the dimensional model, as also studied in [49]. The emotion categories were converted into valence and arousal dimensions following their distribution in Russell’s circumplex model of affect [2]. For valence, the labels were: negative (sad, angry, fearful, disgust), neutral (neutral, surprised) and positive (happy, calm). For arousal, the labels were: low (calm, sad), neutral (neutral, disgust) and high (happy, angry, fearful, surprised).
V Methods
@@
Three approaches were implemented to create the SER models. First,a baseline was established using a standard feature set with classical machine learning algorithms (Section V-A). Second, pre-trained models were used as feature extractors, followed by a linear layer for classification (Section V-B). Finally, a combination of pre-trained models and standard features was analyzed (Section V-C). The Python implementation is available on the following GitHub repository: https://github.com/LuciaGomZa/SER_EMOVOME.git.
For evaluation, the datasets were split into 80% for training and 20% for testing using a speaker-independent approach. To facilitate reproducibility and comparison of results, for testing we used: a proposed set of 20 participants (50% female) available in the EMOVOME Zenodo repository indicated in Section III; session 5 for IEMOCAP; and “fold 0” proposed in [50] for RAVDESS. Details of the label distribution are provided in the Appendix. Additionally, parameter tuning also employed a SI cross-validation scheme with 4 folds for IEMOCAP and 5 folds for EMOVOME and RAVDESS. All methods were evaluated using weighted and unweighted accuracy, described in Section II-B.
V-A Baseline: eGeMAPS and machine learning
@@
The voice messages in EMOVOME were recorded in real-world conditions, using various devices -mostly smartphone microphones- which resulted in different sampling rates (94% at 48 kHz and 6% at 44.1 kHz). To standardize, EMOVOME audio samples were resampled to 44.1 kHz, while RAVDESS and IEMOCAP maintained 48 kHz and 16 kHz, respectively. Next, the extended Geneva Minimalistic Acoustic Parameter Set (eGeMAPS) [51] was extracted with the openSMILE toolkit [52]. The use of this standard set of features facilitates the understanding and reproducibility of the results. The 88 features per audio were normalized by subtracting the mean and dividing by the standard deviation of the development samples. For feature selection, high-correlated features ($p>0.95$) were first eliminated using Pearson’s correlation matrix, and a filter method was then used to select 25%, 50% or 75% of the features based on the highest ANOVA F-values. Finally, Support Vector Machine (SVM) and K-Nearest Neighbours (KNN) models were fitted on the development set according to the cross-validation scheme indicated above. Hyperparameter tuning was performed for both models. For SVM, we changed the kernel (radial basis function, sigmoid), gamma (0.001, 0.01, 0.1, 1, ’auto’, ’scale’) and C (1, 10, 100, 1000). For KNN, we optimized the number of neighbors (1, 3, 5, 7), weights (uniform, distance) and metric (Euclidean, Manhattan, Minkowski). The chosen combination of features and hyperparameters was used to train a model on the entire development set, followed by an evaluation of the test set.
V-B Pre-trained model embeddings
@@
We evaluated several pre-trained models with different architectures, pre-training methodologies and pre-training audio data. First, several variations of the widely used Wav2vec 2.0 model were selected. Considering that EMOVOME is in Spanish, and we wanted to compare it with other two datasets in English, we selected models pre-trained using a multilingual approach: facebook/wav2vec2-xls-r-300m (w2v2-xlsr-128) [53] pre-trained on 436k hours of audios in 128 languages, and facebook/wav2vec2-large-xlsr-53 (w2v2-xlsr-53) [54], pre-trained on 56k hours of audios in 53 languages. Both models included Spanish in the pre-training data. Despite prior research suggesting that fine-tuning models for automatic speech recognition (ASR) do not help with speech emotion recognition [55, 56], we decided to include a fine-tuned model for Spanish ASR. This decision aimed to explore whether this approach outperforms the utilization of a pre-trained model in a different language since [55] and [56] used pre-trained models in the same language as the dataset tested (English). Therefore, we also used the fine-tuned version facebook/wav2vec2-large-xlsr-53-spanish (w2v2-xlsr-53-spa) [54] to obtain the embeddings. We also included a model pre-trained using noisy audios, the model facebook/wav2vec2-large-robust (w2v2-L-robust) [57], as it may be useful for the EMOVOME dataset despite being in English. As a widely-used alternative to Wav2Vec2, we used a HuBERT model, particularly the large version facebook/hubert-large-ll60k (hubert-L) [58]. Additionally, recent studies [41, 59, 60] have indicated that including information about the speaker is helpful for speech emotion recognition. Therefore, following these investigations, two more models were selected: Microsoft’s UniSpeech-SAT-Large (unispeech-L) [61] and a Statistics Pooling Time Delay Neural Network to obtain x-vector embeddings [62]. The former is a model pre-trained using multitask learning, including also the speaker identity during training. The latter provides a speaker embedding learning during a speaker verification task. All the pre-trained models are available in HuggingFace. The x-vector also requires the Speechbrain toolkit [63]. All of them require the input audio to be resampled to 16 kHz.
@@
V-C Pre-trained model embeddings and eGeMAPS